### v1

# Object-X: Learning to Reconstruct Multi-Modal 3D Object Representations

Abstract
Learning effective multi-modal 3D representations of objects is essential for numerous applications, such as augmented reality and robotics. Existing methods often rely on task-specific embeddings that are tailored either for semantic understanding or geometric reconstruction. As a result, these embeddings typically cannot be decoded into explicit geometry and simultaneously reused across tasks. In this paper, we propose Object-X, a versatile multi-modal object representation framework capable of encoding rich object embeddings (e.g., images, point cloud, text) and decoding them back into detailed geometric and visual reconstructions. Object-X operates by geometrically grounding the captured modalities in a 3D voxel grid and learning an unstructured embedding fusing the information from the voxels with the object attributes. The learned embedding enables 3D Gaussian Splatting-based object reconstruction, while also supporting a range of downstream tasks, including scene alignment, single-image 3D object reconstruction, and localization. Evaluations on two challenging real-world datasets demonstrate that Object-X produces high-fidelity novel-view synthesis comparable to standard 3D Gaussian Splatting, while significantly improving geometric accuracy. Moreover, Object-X achieves competitive performance with specialized methods in scene alignment and localization. Critically, our object-centric descriptors require 3-4 orders of magnitude less storage compared to traditional image- or point cloud-based approaches, establishing Object-X as a scalable and highly practical solution for multi-modal 3D scene representation. The code is available at https://github.com/gaiadilorenzo/object-x.
Figure 1: Object-X learns object-centric embeddings from an input object segmentation of a 3D scene reconstruction. The embeddings learned from multi-modal data (e.g., mesh, images, text descriptions) enable fast 3D Gaussian Splat reconstruction via a specifically trained decoder, and other downstream tasks operating directly in the latent space, such as localization and scene alignment. Object-X allows for representing the scene as a set of object descriptors without having to store storage-heavy representations like point clouds and image databases, while providing similiar functionalities.
1 Introduction
Robust 3D scene understanding, incorporating geometric, visual, and semantic information, forms a cornerstone for advances in robotics, augmented reality (AR), and autonomous systems [15, 10, 2]. A key goal is to develop 3D representations that are not only accurate but also compact, efficient, and flexible enough to integrate multiple sensor modalities and support diverse downstream tasks.
Traditional 3D representations, often relying on explicit geometry such as dense point clouds [4, 18] or meshes [16], alongside collections of images, tend to incur prohibitive storage and computational costs. More recently, implicit neural and Gaussian representations, notably Neural Radiance Fields (NeRF) [14] and 3D Gaussian Splatting (3DGS) [9], have achieved state-of-the-art results in synthesising novel views from images, jointly encoding geometry and appearance. However, these methods typically generate a monolithic, scene-level representation, primarily driven by visual input. As a consequence, they inherently lack object-level modularity, making it difficult to reason about individual objects, efficiently incorporate other modalities (e.g., text, semantics), or easily use the representation for object-level tasks beyond rendering or 3D reconstruction.
To address the need for modularity, object-centric approaches, such as 3D scene graphs [1], have gained traction. Such methods decompose a scene into a collection of objects and their relationships, often associating a learned embedding with each object. Such embeddings have proven effective for abstract, object-level tasks, including cross-modal localization [13], scene retrieval [22], and 3D scene alignment [21]. However, a critical limitation persists: existing object embeddings are generally learned for specific tasks and cannot be decoded to reconstruct the explicit, high-fidelity appearance and geometry of the object they represent. This forces systems to retain the original, high-bandwidth source data (images, point clouds, meshes) alongside the learned embeddings, undermining the goals of creating a compact, self-contained, and truly versatile object representation.
In this work, we bridge this crucial gap by introducing Object-X, a framework for learning rich, multi-modal, object-centric embeddings that are simultaneously suitable for downstream tasks and decodable into explicit, high-quality 3D representations. Object-X geometrically grounds multiple input modalities pertaining to an object within a 3D voxel structure, fusing this with semantic attributes (like class labels or object descriptions) to learn a compact, latent embedding. Crucially, we design a decoder that uses this embedding to predict the parameters of a set of 3D Gaussians, enabling high-fidelity, object-level rendering and geometry extraction via 3D Gaussian Splatting. The same learned embedding can be directly leveraged for diverse downstream tasks.
Our experiments on challenging, real-world datasets demonstrate that Object-X supports high-fidelity novel-view synthesis comparable to, and geometric reconstructions superior to, standard 3DGS, while also achieving competitive performance on scene alignment and localization tasks. Critically, Object-X reduces storage requirements by 3-4 orders of magnitude compared to storing the underlying point clouds or images, while offering similar functionalities. Our main contributions are:
  1. 1.
    A novel framework for learning compact, multi-modal, object-centric embeddings that can be decoded into high-fidelity geometry and appearance, parameterized by 3D Gaussians.
  2. 2.
    The demonstration that a single, unified embedding supports both high-quality 3D reconstruction (encompassing novel view synthesis and detailed geometry) and performs competitively on diverse downstream tasks, such as scene alignment and visual localization.
  3. 3.
    Significant storage reduction (3-4 orders of magnitude) compared to explicit representations as the decodability of our embeddings obviates the need to store raw images or point clouds.
Figure 2: Overview of Object-X, learning object embeddings to reconstruct 3D Gaussians and support other tasks such as visual localization [13]. (a) The method takes a mesh or point cloud of an object along with posed images observing it. The canonical object space is voxelized based on object geometry, and DINOv2 features extracted from the images are assigned to each voxel. This produces a $64^{3}\times 8$ structured latent (SLat) representation [29]. (b) The SLat is further compressed into a $16^{3}\times 8$ U-3DGS embedding using a 3D U-Net. The embedding is trained with a masked mean squared error loss to ensure accurate reconstruction of the SLat, which in turn enables decoding into 3D Gaussians using standard photometric losses. (c) Additional task-specific losses, such as those for visual localization [13], can be incorporated to optimize the embedding for multiple objectives.
2 Related Works
Understanding 3D scenes is a fundamental problem in computer vision, with applications spanning robotics, augmented reality, and 3D content creation. Numerous representations have been proposed to capture the complexities of 3D environments, each offering different trade-offs between accuracy, efficiency, storage, and ease of use. Our work, Object-X, builds on and connects several key areas by proposing a novel object-centric embedding strategy.
3D Representations. Traditional methods for representing environments include point clouds [4], meshes [18], and voxel grids [16, 34, 31]. Point clouds offer a direct representation of geometry but lack structure and demand significant storage for detailed scenes. Meshes provide structure by connecting points with polygons, facilitating efficient rendering and geometric operations, though the meshing process can lead to detail loss, especially when optimizing for storage. Voxel grids discretize 3D space, but suffer from high memory requirements at resolutions needed for fine detail. These representations are often accompanied by posed image datasets for applications such as visual localization, which introduces the overhead of managing and storing large image collections. The pursuit of compact and versatile representations motivates our work on learnable object embeddings.
Recently, neural implicit representations have gained attention. Neural Radiance Fields (NeRF) [14] represent scenes with MLPs, enabling photorealistic novel view synthesis but can be computationally intensive. 3D Gaussian Splatting (3DGS) [9] presents a compelling alternative, using a collection of 3D Gaussians for high-quality, real-time rendering. While methods like MV-Splat [3], DepthSplat [30], NoPoSplat [32], and alternatives learn a monolithic 3DGS model for an entire scene from images, we focus on an object-centric paradigm. We learn compact, multi-modal embeddings for individual objects, which can then be decoded into object-level 3DGS parameters. Although recent works have explored 3DGS modifications for editing [33] or compression [23], they typically operate on or refine an existing 3DGS scene. In contrast, Object-X learns fundamental object embeddings that serve not only as a source for 3DGS reconstruction but also as versatile descriptors for various other downstream tasks, offering a more holistic object representation.
3D Scene Graphs. Scene graphs [8] provide a structured representation by capturing objects, their attributes, and interrelations. 3D scene graphs [1] extend this concept to 3D, integrating semantics with spatial and camera information. They have proven useful for tasks like scene alignment [21], retrieval [13, 22], and task planning [5]. The construction of 3D scene graphs has been streamlined by recent advances in object detection and relationship prediction, with tools such as MAP-ADAPT [34], OpenMask3D [24], and ConceptGraphs [5]. Existing scene graph methodologies often associate nodes with learned embeddings tailored for specific tasks (e.g., alignment, retrieval). However, a key limitation is that these embeddings typically lack a generative or reconstructive capability; they cannot be decoded back into explicit object geometry or appearance. This necessitates retaining the original sensor data (images, point clouds) alongside the graph, undermining compactness. Object-X addresses this gap by learning rich, multi-modal object embeddings that are explicitly designed to be decodable into high-fidelity 3DGS representations. Furthermore, these same embeddings retain strong descriptive power, enabling competitive performance on downstream tasks like localization and alignment without requiring task-specific modifications. This dual capability – reconstructive and descriptive – is a core contribution of our work, offering a pathway to leverage the semantic richness while also providing access to explicit 3D object representations.
3D Generative Methods for content creation have rapidly advanced, with 3DGS [9] emerging as an expressive and efficient primitive. Several methods leverage 2D distillation from text or images to generate 3D 3DGS scenes [25, 19]. More recent works explore structured latent spaces for improved scalability and control in generation. For instance, Trellis [29] uses a unified Structured LATent (SLat) representation, decodable into various 3D forms including Gaussians, for large-scale object generation. L3DG [20] employs a latent diffusion framework with VAEs for efficient sampling and 3DGS rendering, while DiffGS [35] introduces a diffusion model for controlling Gaussian parameters.
These methods focus on de novo generation or generation from abstract inputs (e.g., text prompts, style images), showcasing the potential of learned latent spaces for 3D content creation. Object-X shares the goal of leveraging learned latents but differs in its main objective. Instead of open-ended generation, our focus is on learning compact and versatile embeddings from observed multi-modal data. The key is to create embeddings that capture an object geometry and appearance for reconstruction, while also being discriminative enough for tasks like localization and alignment. Thus, we emphasize robust representation learning from captured data for multi-task utility, rather than pure synthesis.
3 Learning Versatile Object Embeddings
We propose Object-X, taking a reconstructed scene with a 3D object segmentation as input and learning a compact and descriptive embedding for each object from their associated multi-modal data (e.g., images, point cloud). To achieve this, we process each 3D object through the following steps:
  1. 1)
    Extract Structured Latent Representation (SLat): We first process the input data for an object. We voxelize the object into a canonical 3D grid and aggregate local image features within these voxels via multi-view projection, inspired by [29]. A 3D encoder then transforms these voxel-aligned features into a structured set of latent vectors (the SLat), which represents the object’s initial rich encoding.
  2. 2)
    Project SLat to an Unstructured Embedding: We then learn to compress SLat into a dense and unstructured latent representation of a fixed, significantly smaller dimension. Besides compression loss, other objectives are also considered during training to facilitate downstream tasks, e.g., object retrieval. This embedding, which we term the U-3DGS Embedding [^1] Named for its designed decodability into 3D Gaussian Splats, though it serves broader purposes., becomes our core storage unit and descriptor for an object.
  3. 3)
    Decode the U-3DGS Embedding to 3DGS: One key application of the U-3DGS embedding is its decodability. For reconstruction, we map this dense embedding back to an SLat representation and then decode this into a full 3D Gaussian Splat model for the object.
Below, we detail the formation of the SLat representation. We then describe its compression into the versatile U-3DGS embedding. Following this, we explain the decoding mechanism that enables 3DGS-based reconstruction from this embedding. The utility of U-3DGS for other direct downstream applications will be demonstrated in subsequent sections. Fig. 2 summarizes the overall pipeline.
3.1 Structured Latents from Multi-View Images
Let a set of object instances $\mathcal{O}$ be given, where each object $o=(\mathcal{P},\mathcal{I},\mathcal{M},\mathcal{A},\dots)\in\mathcal{O}$ is associated with a multi-modal input, such as multi-view images ($\mathcal{I}$), instance masks ($\mathcal{M}$), point clouds ($\mathcal{P}$) and other attributes ($\mathcal{A}$). For each object $o$, we extract an object-specific latent embedding $\mathbf{w}_{o}$. The scene representation is defined as $\mathcal{W}=\{\mathbf{w}_{o}\}_{o\in\mathcal{O}}$, enabling lightweight storage compared to dense point clouds or images. The goal is to learn mapping $\mathcal{O}\;\mapsto\;\mathcal{W}$, where $\mathcal{W}$ allows decoding into 3D objects. Here, we focus on modalities $\mathcal{P}$ and $\mathcal{I}$ to learn the reconstructable embedding. Other modalities will be discussed when training also for downstream applications in the next section.
Voxelization and Feature Extraction. Given multi-view images of an object $o$, we first voxelize its canonical 3D space into an $N\times N\times N$ grid (e.g., $64^{3}$). For each voxel $p_{i}$ that intersects the surface of the object, we project $p_{i}$ into each image to retrieve localized features. Similar to [29], we employ a pre-trained DINOv2 feature encoder [17] on masked object images. Averaging these image-level features yields a per-voxel feature $f_{i}\in\mathbb{R}^{D}$. Concatenating over all active voxels forms a sparse, voxel-aligned feature set as $f\;=\;\{(f_{i},p_{i})\}_{i=1}^{L},L\ll N^{3}$.
Structured Latent Representation (SLat). We convert $f$ into a structured latent representation $z=\{(z_{i},p_{i})\}_{i=1}^{L}$ via a 3D encoder $E$ as follows:
$z=\{(z_{i},p_{i})\}_{i=1}^{L}\;=\;\mathcal{E}(f),\;z_{i}\in\mathbb{R}^{C},\;p_{i}\in\{0,\dots,N-1\}^{3}.$
Each voxel $p_{i}$ is paired with a local latent $z_{i}$ capturing shape and appearance. By preserving the sparse grid structure, $z$ offers geometric grounding (through $p_{i}$) and localized feature encoding (through $z_{i}$).
3.2 Compression to a Dense Embedding
Although $z$ is sparse relative to raw 3D data, it can still be large for high-resolution grids when we aim at capturing fine details. Moreover, downstream tasks favor a compact embedding for each object, rather than a collection of per-voxel vectors. Thus, we learn a mapping to compress $z$ into a fixed-size vector $\mathbf{w}\in\mathbb{R}^{d}$, where $d$ is small and independent of the size of the object at hand as $\mathbf{w}=f_{\text{U-3DGS}}(z)$, where $\mathbf{w}$ is an Unstructured 3D Gaussian Splatting (U-3DGS) embedding. We implement $f_{\text{U-3DGS}}$ using a 3D network (similar to a U-Net) that first organizes $z$ into a dense $N\times N\times N\times C$ tensor (with zero-filling for inactive voxels), then, downsamples and encodes it to $\mathbf{w}$.
Masked MSE for Compression Learning. We supervise $f_{\text{U-3DGS}}$ by requiring that the dense embedding $\mathbf{w}$ can be decoded back into the original structured representation $z$ (or a close approximation). Specifically, we introduce a decompression function $f_{\text{decomp}}$ that maps $\mathbf{w}$ back to a predicted $\hat{z}=\{\hat{z}_{i},p_{i}\}_{i=1}^{L}$. We then encourage $\hat{z}_{i}\approx z_{i}$ under a masked mean-square-error (MSE) or L1 loss that focuses on occupied voxels as $\hat{z}\;=\;f_{\text{decomp}}(\mathbf{w})$, and the objective includes
$\mathcal{L}_{\text{compress}}\;=\;\frac{1}{L}\sum_{i=1}^{L}\Bigl{[}M_{i}\,\lVert\hat{z}_{i}-z_{i}\rVert^{2}\;+\;\tfrac{1-M_{i}}{w}\lVert\hat{z}_{i}-z_{i}\rVert^{2}\Bigr{]},$
where binary mask $M_{i}$ indicates if voxel $p_{i}$ is occupied, and $w$ is a down-weighting factor for non-occupied regions.
3.3 Decoding to 3D Gaussian Splats
In this section, we use the learned object embeddings $\mathbf{w}$ to reconstruct the object into a 3DGS representation. As discussed in the previous section, we can obtain the reconstructed structured latent representation $\hat{z}$ from $\mathbf{w}$ by applying decompression network $\hat{z}\;=\;f_{\text{decomp}}(\mathbf{w})$.
Deterministic Autoencoder for 3DGS. We next train a decoder $D_{\text{GS}}$ that takes in the reconstructed structured latents $\hat{z}$ (obtained from $\mathbf{w}$) and outputs a set of 3D Gaussian Splat parameters as follows:
$\Theta\;=\;\mathcal{D}_{\text{GS}}(\hat{z}),\quad\Theta\;=\;\Bigl{\{}\!(x_{i},s_{i},q_{i},\alpha_{i},c_{i})\!\Bigr{\}}_{i=1}^{N}.$ (1)
Each Gaussian is specified by position $x_{i}$, scale $s_{i}$, rotation $q_{i}$, opacity $\alpha_{i}$, and color $c_{i}$. We train $D_{\text{GS}}$ by rendering these Gaussians from multiple viewpoints and minimizing image-space reconstruction losses (e.g., L1, SSIM, and LPIPS) against ground-truth images of the object using loss $\mathcal{L}_{\text{render}}\;=\;\lambda\,\mathcal{L}_{\text{L1}}\;+\;(1-\lambda)\bigl{[}1-\mathcal{L}_{\text{SSIM}}\bigr{]}\;+\;\mathcal{L}_{\text{LPIPS}}$. Since $\hat{z}$ can accurately encode fine object details, $D_{\text{GS}}$ learns to produce high-fidelity splats.
Voxel-Level Offsets. For spatial alignment, the center of each Gaussian is computed as $x_{i}=p_{i}+\tanh\bigl{(}o_{i}\bigr{)}$, where $o_{i}$ is an offset predicted by $D_{\text{GS}}$ and $p_{i}$ is the voxel location. This ensures positions remain near the coarse voxel layout, but can adjust locally for more precise fits.
Training. The proposed pipeline is end-to-end trainable, learning a compact embedding space in a single training pass. The optimization jointly minimizes the reconstruction loss, ensuring accurate recovery of the SLat representation from the U-3DGS embedding, and photometric loss, learning the mapping from SLat features into 3DGS that can be used for NVS and surface reconstruction.
Figure 3: The proposed Object-X learns per-object embeddings that are beneficial for a number of downstream tasks, besides object-wise 3DGS reconstruction, such as cross-modal visual localization [13] (via image-to-object matching), 3D scene aligment [21] (via object-to-object matching), and full-scene reconstruction by integrating per-object Gaussians primitives.
3.4 Learning Auxiliary Tasks
The U-3DGS embedding, primarily learned to capture object appearance and geometry for high-fidelity reconstruction, also serves as a potent foundation for various downstream auxiliary tasks such as visual localization [13] and 3D scene alignment [21, 22] (see Fig. 3). To further enhance performance on such tasks, we augment the pre-trained U-3DGS representation by integrating information from other relevant modalities. This augmentation involves incorporating features from auxiliary data sources like textual descriptions, object relationships, or broader scene context (which can be derived, for example, from a 3D scene graph structure [21]). Each auxiliary modality is processed by its own dedicated encoder (e.g., a CLIP-based model with a projection head for text features), following [21]. The resulting feature vectors from these auxiliary encoders are then concatenated with the original U-3DGS embedding to form a richer, multi-faceted representation for the object. The training strategy for these augmented representations, leveraging the pre-trained U-3DGS encoder and decoder, proceeds in two main stages after the initial U-3DGS pre-training:
Auxiliary Encoder Training with Frozen Core: Initially, the pre-trained U-3DGS encoder and decoder (responsible for appearance and geometry) are kept frozen. The newly introduced encoders for the auxiliary modalities are trained. In this stage, the learning is guided by a combined loss:
$\mathcal{L}_{\text{aux}}=\mathcal{L}_{\text{task}}(\mathbf{w}_{\text{concat}})+\lambda_{\text{recon}}\mathcal{L}_{\text{recon}}(\mathbf{w}_{\text{U-3DGS}}),$ (2)
where $\mathbf{w}_{\text{concat}}$ is the full concatenated embedding (U-3DGS + auxiliary features). $\mathcal{L}_{\text{task}}$ is the task-specific loss (e.g., a contrastive loss for localization). $\mathcal{L}_{\text{recon}}$ is the reconstruction loss also used for pre-training the U-3DGS embeddings. Crucially, $\mathcal{L}_{\text{recon}}$ is applied using the U-3DGS decoder, which operates only on the corresponding U-3DGS portion ($\mathbf{w}_{\text{U-3DGS}}$) of $\mathbf{w}_{\text{concat}}$. This ensures that the new auxiliary features are learned in a way that remains compatible with, and does not corrupt, the reconstructive capabilities of the core U-3DGS representation. $\lambda_{\text{recon}}$ balances these two objectives.
Joint Fine-tuning: After the auxiliary encoders have been trained, all network components are unfrozen. This includes the U-3DGS encoder and decoder, as well as all auxiliary modality encoders. The entire ensemble is then fine-tuned end-to-end using the same combined loss $\mathcal{L}_{\text{aux}}$ (Eq 2).
Through this process, Object-X learns to effectively fuse intrinsic object properties (geometry, appearance via U-3DGS) with extrinsic or contextual information (text, relationships via auxiliary encoders). We will demonstrate this approach by training our model for visual localization [13], and subsequently evaluate its zero-shot or fine-tuned performance on related tasks like 3D scene alignment, showcasing the versatility and robustness of the learned augmented embeddings.
Figure 4: Object reconstructions. Each row shows an input object (left) and its reconstruction obtained by, from left to right: (i) 3DGS [27] optimized on all images, (ii) 3DGS or (iii) 2DGS [37] using only 12 multi-view images, and (iv) Object-X. For each method, we present a rendered image from the reconstructed 3D Gaussians and the corresponding mesh.
4 Experiments
Next, we will provide experiments on various tasks benefiting from Object-X. Ablation studies, more visuals, and detailed descriptions of baselines are provided in the supplementary material.
Mesh extraction. To evaluate the geometric accuracy, we extract a triangle mesh from the optimized 3D Gaussians, following the procedure from 2DGS [37]. We first render depth maps from different viewpoints. These depths are fused using the Truncated Signed Distance Function (TSDF) integration in Open3D [36]. Finally, a triangle mesh is extracted using Marching Cubes [12].
Implementation details. All experiments are conducted on a machine with an A100 GPU with 80GB of RAM. During sparsification, a threshold of 0.5 is applied to the predicted occupancy. The mesh is constructed using a voxel size of 0.015 and an SDF truncation value of 0.04.
Datasets. The 3RScan dataset [26] consists of 1,335 annotated indoor scenes covering 432 distinct spaces, with 1,178 scenes (385 rooms) used for training and 157 scenes (47 rooms) reserved for validation and testing. The dataset provides semantically annotated 3D point clouds, with certain scenes captured over extended periods to reflect environmental changes. Scene graph annotations are available from [27]. Since the test set lacks such annotations, we reorganized the original validation split, allocating 34 scenes (17 rooms) for validation and 123 scenes (30 rooms) for testing. Objects without available images were removed to ensure a consistent evaluation.
ScanNet. To evaluate generalization, we test on ScanNet [4] without training our model on it. Since ScanNet does not provide scene graph annotations, we apply SceneGraphFusion [28] on RGB-D sequences to generate 3D instance segmentations and object relationships (used for auxiliary tasks). Compared to 3RScan, ScanNet captures RGB-D sequences at a higher frame rate with minimal motion between consecutive frames. To ensure diverse viewpoints, we sample one image every 25 frames. We use 77 test scenes from the split defined in [13]. Scenes, where SceneGraphFusion fails to generate annotations, are excluded. As in 3RScan, objects without associated images are discarded. The test split, along with its annotations, will be publicly released.
Table 1: 3DGS Object reconstruction photometric quality, geometric accuracy, runtime, and storage efficiency on 3RScan [26] and ScanNet [4]. We compare Object-X with baselines that store objects as a set of 12 (3RScan) or 4 images (ScanNet) and reconstruct 3D Gaussians at test time using 3DGS [27], 2DGS [37], and DepthSplat [30]. As a reference, we report the results of 3DGS, optimizing directly on all dataset images. We report NVS scores (SSIM, PSNR, LPIPS), geometric accuracy (Accuracy, Completion, and F1 score at a 0.05 m threshold), per-object run-time (secs), and storage (MB). We do not show geometric accuracy for DepthSplat as it failed mesh reconstruction.
Method SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ Acc.@0.05 $\uparrow$ Compl.@0.05 $\uparrow$ F1@0.05 $\uparrow$ Time (s) $\downarrow$ Storage (MB)
3RScan 3DGS [27] 0.956 34.009 0.051 33.75 73.38 41.41 58.461 32.74
[…]
2DGS (12V) [7] 0.932 29.613 0.093 26.55 67.69 35.67 84.280 16.71
DepthSplat (12V) [30] 0.619 21.669 0.304 - - - 0.491 16.71
Object-X 0.953 30.981 0.065 80.22 77.80 77.80 0.051 10.14
ScanNet 3DGS [27] 0.975 38.138 0.032 33.81 79.13 43.00 129.02 270.97
[…]
2DGS (4V) [7] 0.945 29.604 0.073 26.96 74.85 37.10 169.65 155.26
DepthSplat (4V) [30] 0.832 26.152 0.134 - - - 0.138 155.26
Object-X 0.966 31.563 0.047 88.66 90.08 89.09 0.033 100.14
1. Object Reconstruction. First, we evaluate the Object-X decoder in terms of storage efficiency, geometric fidelity, and visual quality on the object reconstruction task.
Baselines. 3DGS [9] serves as a high-fidelity baseline, representing each object as a set of 3D Gaussians. While this approach captures fine details, it requires substantial storage, as every object is represented by a set of Gaussians. We provide results for 3DGS (12 views) that stores each object as 12 images captured from different viewpoints. During reconstruction, 3DGS is applied to recover the 3D Gaussians from these views. This reduces storage compared to full 3DGS but introduces a trade-off: reconstruction takes longer, and the quality may be slightly degraded. 2DGS (12 views) [7] follows the same 12-image storage strategy but employs 2DGS [37] instead of 3DGS. Both 2DGS and 3DGS leverage the segmented mesh as initialization. Default parameters are used. DepthSplat (12 views) [30] also relies on 12 stored views but reconstructs objects using DepthSplat, a fast feed-forward network. We use the pre-trained model provided by the authors, which was not trained on 3RScan. Since we train on 3RScan, comparisons on this dataset may be unfavorable to DepthSplat. However, we also evaluate on ScanNet, where neither DepthSplat nor our method has been trained.
For the baselines, we select $k$ frames that maximize viewpoint diversity. To achieve this, we apply $k$-means clustering to the positions of the cameras observing the object. In 3RScan, we use the maximum number of frames supported by DepthSplat (i.e., 12), while in ScanNet, where fewer non-overlapping frames are available, we limit the selection to four frames per object.
We present examples in Fig. 6, showing renderings and the reconstructed meshes. Object-X produces significantly smoother renderings and higher-quality meshes, whereas meshes reconstructed by baselines exhibit strong artifacts and fail to achieve accurate geometry.
Metrics. We evaluate our method using the standard novel view synthesis scores: PSNR, SSIM, and LPIPS. Additionally, we report standard geometric metrics: accuracy, completeness, and F1 score.
Results on 3RScan. The top part of Table 1 reports the results on 3RScan. Object-X achieves comparable novel view synthesis quality to the baselines, with an SSIM score closest to 3DGS, the second-best PSNR score, and the best LPIPS score. Object-X substantially outperforms all baselines in geometric accuracy. Our method improves geometric accuracy by a large margin of 46 percentage points compared to all baselines while also exhibiting good completeness. This demonstrates that the proposed U-3DGS embeddings effectively capture object geometry, accurately recovered by the Object-X decoder. We omit geometric results for DepthSplat [30], which failed to produce reasonable geometry. Our runtime is three orders of magnitude faster than methods relying on optimization. Moreover, our approach requires an order-of-magnitude less storage, as we only store a single embedding per object instead of numerous images or 3D Gaussians.
Results on ScanNet. The bottom part of Table 1 reports the results on ScanNet. Note that we did not train our model on this dataset and used the model trained on 3RScan. Even without training, we achieve the highest novel view synthesis scores compared to the baselines, being the closest to the reference 3DGS reconstruction. Our geometric accuracy significantly outperforms all baselines. Given the higher resolution images in this dataset compared to 3RScan, the optimization-based approaches are quite slow, even when using only 4 views. We are four orders of magnitude faster than 3DGS (4V) and 2DGS (4V) requiring $3$ ms to reconstruct an object on average.
2. Scene Reconstruction. Although the we do not explicitly perform scene reconstruction, we evaluate scene composition by integrating object-level reconstructions. Object-X composes the full scene by decoding embeddings for each object and rendering their splats jointly. While this method is highly efficient, artifacts can emerge due to missing object segmentations. We also show results for Object-X + Opt, where the initial scene is constructed via Object-X, and optimized using 3DGS on the same image set used in 3DGS (12V). This setup isolates the benefit of Object-X as a strong initialization while allowing to overcome the problems caused by missing objects.
Table 2 reports results on the 3RScan dataset. While Object-X achieves lower SSIM and PSNR compared to 3DGS (12V), it significantly outperforms all methods in geometric accuracy. Also, it runs two orders of magnitude faster than the baselines. With refinement (Object-X + Opt), our method not only matches or exceeds the photometric quality of 3DGS (12V), but also achieves the highest geometric accuracy by a large margin. Runtime remains comparable to the 12-view baselines, and significantly faster than full-scene 3DGS optimization.
3. Single-Image Reconstruction. We further evaluate the flexibility of Object-X by reconstructing objects from a single RGB or RGB-D frame. Let us note that the Object-X was not explicitly trained for this task nor to infer unseen parts of an object. In the RGB-only scenario, we estimate monodepth using [6]. Using the object mask and depth map, we lift the object pixels in 3D, obtaining a point cloud. We then voxelize it following the process described in Sec. 3.1. Object-X is then applied to obtain the object embedding from this input which is then fed directly into our decoder. We compare our approach to 3DGS [27] which optimizes 3D Gaussian splats based on a single masked image. Results in Table 1a show that Object-X outperforms 3DGS in all metrics on RGB-D and RGB inputs while also achieving significantly better F1 scores, indicating more precise and reliable reconstructions. We note that while there is still room for improvement in this task, achieving improved results demonstrate the versatility of the learned object embeddings.
Table 2: Full-scene composition on 3RScan [26]. We compare Object-X to 3DGS [27] optimized on all unmasked images, and two 12-view baselines: 3DGS (12V) and 2DGS (12V), which optimize scenes using a subset of training images constructed by taking the union of the 12 best views selected per object. Object-X achieves the second highest geometric accuracy the fastest. When combined with refinement (Object-X + Opt), it also achieves the best perceptual quality among all methods.
Method SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ Acc.@0.05 $\uparrow$ Compl.@0.05 $\uparrow$ F1@0.05 $\uparrow$ Time (s) $\downarrow$
3RScan 3DGS [27] 0.855 24.404 0.417 16.48 31.99 21.40 260
[…]
2DGS (12V) [7] 0.752 18.900 0.523 18.09 30.92 22.61 250
Object-X 0.677 15.488 0.526 46.22 54.98 49.99 1
Object-X + Opt 0.727 17.098 0.507 63.00 71.54 66.49 150
Method LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ F1@0.05 $\uparrow$
RGB-D 3DGS 0.131 26.58 0.891 27.01
Object-X 0.099 28.03 0.926 44.79
RGB 3DGS 0.129 26.72 0.896 18.70
Object-X 0.110 27.27 0.918 10.53
(a) Single-Image Object Reconstruction on RGB (depth predicted by [6]) and RGB-D frames from 3RScan [26]. Photometric and geometric accuracy of 3DGS vs. Object-X.
Method Modalities 10 scenes
[…]
SGLoc [13] 53.6 81.9 92.8
CrossOver [22] 46.0 77.9 90.5
Object-X 56.6 82.2 91.8
(b) Coarse Visual Localization on 3RScan [26]. Retrieval recall at various thresholds using various methods and input modalities.
Figure 5: Comparison of (a) object reconstruction and (b) coarse localization performance using 3DGS and Object-X across tasks and input modalities.
4. Coarse Visual Localization. We evaluate visual localization on 3RScan. Following the protocol from SceneGraphLoc [13], we evaluate on 123 scenes from 30 rooms in the test set. We select query images for each scene and match them against 10 candidate scenes (including the target) to determine if the correct scene can be identified. This process is repeated for every image in each room, resulting in a total of 30,462 query images used for evaluation. Our method, along with SceneGraphLoc and the recent CrossOver [22], uses a ViT to extract per-patch object embeddings from the query image. For each candidate scene, a similarity score is calculated by identifying the most similar object embedding in the scene for each patch in the image. Robust voting is then performed across all patch-object similarities to determine the scene where the query image was captured. We report the scene retrieval recall at 1, 3, and 5, measuring the percentage of queries for which the correct scene is ranked within the top 1, 3, and 5 retrieved scenes. The results are shown in Table 1b. We indicate whether a method uses point cloud ($\mathcal{P}$), image ($\mathcal{I}$), other modalities like object attribute and relationship ($\mathcal{O}$), or the proposed U-3DGS embedding. Our method, leveraging 3DGS and other modalities, achieves the highest R@1 and R@3 scores.
5. 3D Scene Alignment. To further assess our generalization capabilities, we evaluate on a new task, Scene Alignment, on 3RScan, following the protocol from SGAligner [21]. This task involves matching objects across partially overlapping scans of the same scene by comparing their embeddings. Unlike SGAligner, explicitly trained for this task using point cloud and object-level modalities, our method relies solely on the proposed Object-X embedding trained with reconstruction and localization losses, without finetuning for scene alignment. As shown in Table 3, Object-X performs comparably in all metrics (achieving the second highest accuracy) to the baselines tailored for this task.
Table 3: 3D Scene Alignment on 3RScan [26] by Object-X, SGAligner [21] and EVA [11]. We report Mean Reciprocal Rank and Hits@K that denotes the proportion of correct matches appearing within the top $K$, based on cosine similarity. Evaluations are conducted using modalities: point cloud ($\mathcal{P}$), others ($\mathcal{O}$), and 3DGS. In constrast to the baselines, Object-X is used without training on this task.
Method Modalities Mean RR $\uparrow$ Hits@1 $\uparrow$ Hits@2 $\uparrow$ Hits@3 $\uparrow$ Hits@4 $\uparrow$ Hits@5 $\uparrow$
$\mathcal{P}$ $\mathcal{O}$ 3DSG
EVA [11] 0.867 0.790 0.884 0.938 0.963 0.977
SGAligner [21] 0.884 0.835 0.886 0.921 0.938 0.951
SGAligner [21] 0.950 0.923 0.957 0.974 0.982 0.987
Object-X 0.910 0.864 0.917 0.948 0.965 0.975
5 Conclusion
We introduced Object-X, a novel framework for learning compact, versatile, multi-modal object-centric embeddings. These unique embeddings are decodable into high-fidelity 3D Gaussian Splats for object reconstruction while also serving as potent descriptors for diverse downstream tasks, such as visual localization, single-image reconstruction, and 3D scene alignment, often without task-specific fine-tuning. Object-X achieves excellent geometric accuracy and novel-view synthesis, comparable or superior to specialized methods, while drastically reducing storage requirements by 3-4 orders of magnitude by obviating the need for raw sensor data. Our approach effectively bridges the gap between abstract learned object representations and detailed explicit 3D models, offering a scalable and practical solution for advanced 3D understanding. The code will be made public.
Limitations. Despite these advances, Object-X has limitations. The high degree of compression can lead to a loss of the finest details in some reconstructions, particularly for large or complex objects. Furthermore, while promising in zero-shot scenarios for tasks like single-image object reconstruction, performance may not yet consistently match optimized task-specific methods.
Supplementary Material
This supplementary material provides additional training details, experimental setups, visualizations, and ablation studies in support of the main paper. It is organized as follows:
  1. 1.
    Additional details on baseline comparisons and visualizations for object-level, image-based, and scene-level reconstruction (Section A)
  2. 2.
    Training procedures for pretraining, compression, and downstream adaptation (Section B)
  3. 3.
    Setup and evaluation details for visual localization using U-3DGS embeddings (Section C)
  4. 4.
    Scene alignment task setup, including sub-scene construction and evaluation metrics (Section D)
  5. 5.
    Ablation results on compression, occlusion robustness, and architectural variants (Section E)
Appendix A Baselines and Visualizations
This section details the implementation and evaluation protocols for all baseline methods discussed in the main paper. We cover object-level, scene-level, and single-image reconstruction settings. To ensure a fair and rigorous comparison across all experiments, we maintain consistent supervision levels, initialization strategies, and evaluation metrics when assessing storage requirements, visual quality, and geometric fidelity.
A.1 Object Reconstruction Baselines
All optimization-based methods, specifically 3D Gaussian Splatting (3DGS) [9] and 2DGS [37], operate on masked RGB-D input sequences. Our comparative analysis includes three primary optimization-based baselines:
  1. 1.
    3DGS (Full Scene): Utilizes all available images from a given scene to serve as an upper-bound reference reconstruction.
  2. 2.
    3DGS ($k$V): Employs $k$ pre-selected views that observe the target object.
  3. 3.
    2DGS ($k$V): Also uses $k$ pre-selected views observing the target object.
To ensure a fair comparison by providing identical starting conditions for all methods, we initialize Gaussian splats directly from ground-truth object meshes. The $k$ views for object-specific baselines are selected using a $k$-means clustering strategy (detailed below) to promote viewpoint diversity and ensure high object visibility. Crucially, evaluation is consistently performed on a disjoint test set of images that were not utilized during the training or optimization phases of any method. For both 3DGS and 2DGS, we conduct optimization for 7,000 iterations using their default hyperparameter settings. For experiments on the 3RScan [26] dataset, we use 12 views and, on ScanNet [4], which typically offers fewer views per object instance, we restrict the number of selected views, $k$, to a maximum of 4 per object.
Regarding DepthSplat [30], we evaluate using the publicly available pre-trained model. As this model was not trained on object reconstruction tasks, we apply it to the unmasked image and, we post-process its output by removing any splats that fall entirely outside the masked object region.
Visual comparisons are provided in Figure 6. Alongside rendered novel views, we present mesh reconstructions derived from the 3D Gaussians using the TSDF fusion technique, as proposed in [37]. These visualizations demonstrate that our proposed method, Object-X, achieves significantly smoother novel view syntheses and more geometrically accurate mesh reconstructions compared to the baselines.
Frame Selection Protocol. To ensure consistent and representative view selection across all relevant experiments (object-level and scene-level $k$-view baselines), we employ a clustering-based strategy for choosing training/optimization views. From the available set of frames for an object or scene, we first cluster their camera extrinsics (position and orientation) using $k$-means. Subsequently, we select one frame from each resulting cluster, prioritizing the frame that exhibits the fewest masked pixels (i.e., maximal object visibility within the frame). Any objects for which no valid test images remain after this selection process (e.g., due to insufficient visibility in all remaining frames not reserved for testing) are excluded from the evaluation set to maintain fairness.
Figure 6: Object reconstructions. Each row shows an input object (left) and its reconstruction obtained by, from left to right: (i) 3DGS [27] optimized on all images, (ii) 3DGS or (iii) 2DGS [37] using only 12 multi-view images, and (iv) Object-X. For each method, we present a rendered image from the reconstructed 3D Gaussians and the corresponding mesh.
A.2 Scene-Level Reconstruction
For full-scene evaluation, we adapt 3DGS and 2DGS to operate jointly across all objects within a scene. This is achieved by utilizing the union of the same $k$ views per object as in the object-level experiments (specifically, 12 views for 3RScan and 4 for ScanNet), but here the RGB-D inputs are unmasked. Our method, Object-X, reconstructs the scene by independently decoding the learned U-3DGS embedding for each constituent object and then rendering their collective splats. This compositional approach requires no additional scene-level optimization. We also evaluate an augmented version, denoted as Object-X + Opt. This variant leverages the compositional scene from Object-X as an initialization for a subsequent refinement stage. Specifically, it undergoes an additional 4,000 iterations of 3DGS optimization. To ensure stability during this fine-tuning process, all learning rates are reduced by a factor of 10×10$\times$10 × compared to the standard 3DGS settings.
Visualizations of scene-level reconstructions are presented in Figure 8. While Object-X generally produces significantly smoother results than the baseline methods, its performance can be affected by objects missing from the input segmentations (e.g., a poster on a wall, as shown in the first row of the figure, or the objects on the desk, as shown in the second row). Additionally, fine-grained details might sometimes be diminished. However, applying 3DGS optimization as a post-processing step (Object-X + Opt) yields substantial improvements in accuracy, effectively recovering such lost details.
A.3 Single-Image Reconstruction
We extend our evaluation to a single-view reconstruction setting for all methods. In this scenario, 3DGS is optimized from scratch using a single masked RGB-D image (and its corresponding RGB image) for 3,000 iterations. To ensure a fair comparison, Object-X utilizes the same reference image. This image is selected based on criteria that maximize unmasked object coverage while minimizing cropping along the image borders. We also test our method with only RGB input with depth predicted by Metric3D [6] to generate the initial point cloud. Similarly to the scene reconstruction case, we use Object-X to provide an initial reconstruction which we further refine by applying an additional 1,000 iterations of 3DGS optimization, using learning rates reduced by a factor of 10×10$\times$10 × (consistent with the scene-level refinement). Visual results for this setting are presented in Figure 7.
Notably, despite Object-X not being explicitly trained for single-image reconstruction tasks, it frequently produces visually cleaner reconstructions than 3DGS when both methods are constrained to the same single input view and 3DGS is optimized from scratch under these conditions. The reconstructed meshes are also substantially more accurate than the ones from 3DGS.
A.4 Evaluation Summary
Across all experimental settings, methods are evaluated on a fixed set of test views, distinct from training/optimization views, on a per-object or per-scene basis as appropriate. We report standard quantitative metrics, Peak Signal-to-Noise Ratio (PSNR), and the perceptual metric LPIPS. Qualitative comparisons are provided in the relevant figures accompanying each experimental section ( e.g., Figure 6 for object-level, Figure 8 for scene-level, and Figure 7 for single-image results). Across all evaluated levels – single-view, multi-view object reconstruction, and full-scene composition – Object-X demonstrates strong performance, simultaneously offering significant advantages in terms of computational efficiency and flexibility in initialization.
Figure 7: Qualitative comparison for image to 3D. We compare the proposed Object-X to standard 3DGS [27] on RGB and RGB-D inputs. For each method, we present the image from which the object (left column) is reconstructed and the rendered novel view together with the mesh reconstructed from the 3D Gaussians by: (2nd column) proposed Object-X with RGB-D input; (3rd) 3DGS with RGB-D; (4th) proposed Object-X with RGB input; (5th) 3DGS with RGB. The proposed method leads to significantly cleaner novel views and meshes than 3DGS applied to a single image.
Figure 8: Qualitative comparison for full-scene composition. We compare the proposed Object-X to standard 3DGS [27] optimized on all unmasked scene images, and two 12-view baselines: 3DGS (12V) and 2DGS (12V), which optimize scenes using a subset of training images constructed by taking the union of the 12 best views selected per object.
Appendix B Training Details
The training procedure encompasses three primary phases: sparse representation learning, compression model training, and adaptation for downstream tasks. Each phase employs distinct optimization settings to ensure both stability and efficiency. For the sparse transformer-based encoder and decoder, we apply gradient clipping at a threshold of $0.01$. This is crucial for stabilizing the training process and preventing excessively large updates within the structured latent space. Optimization is conducted using the AdamW optimizer with a learning rate of $1\times 10^{-4}$. This learning rate is selected to strike an effective balance between training stability and convergence speed. AdamW is chosen for its decoupled weight decay mechanism, which aids in regularizing the model without adversely affecting the gradient-based optimization updates. During the compression phase, a 3D U-Net architecture is trained to map the structured latent representation to a more compact form suitable for efficient storage or transmission. A higher learning rate of $1\times 10^{-3}$ is utilized in this phase. This facilitates accelerated convergence while preserving reconstruction quality. Explicit gradient clipping is not deemed necessary for the U-Net, as its inherent hierarchical structure and typical training dynamics provide sufficient stabilization.
For adaptation to downstream tasks, such as object localization or instance retrieval, training is performed using the AdamW optimizer with a learning rate of $1\times 10^{-3}$ when the voxel-based latent representation is kept frozen. However, if the voxel representation is fine-tuned concurrently with the task-specific modules, a lower learning rate of $1\times 10^{-4}$ is adopted. This approach helps to mitigate the risk of catastrophic forgetting of the learned representations. Key regularization techniques employed include the aforementioned gradient clipping and structured weight decay (e.g., as provided by the AdamW optimizer).
Appendix C Supplementary: Coarse Visual Localization
We provide additional details for the visual localization experiment on the 3RScan dataset. This experiment is designed to evaluate the downstream utility of the U-3DGS embeddings when augmented with auxiliary modalities.
Training. The model utilized for localization is trained following the auxiliary learning setup described in the main paper. We freeze the pre-trained U-3DGS encoder and decoder. Auxiliary encoders are then trained on object-level inputs derived from 3D scene graphs, specifically object relationships, attributes, and structural context. Each RGB query image is processed through a DINOv2 backbone followed by a patch-level encoder to generate patch-wise descriptors. A contrastive loss function aligns these image patches with the corresponding object embeddings within the scene. Concurrently, a compression loss ensures that the original U-3DGS component of the joint embedding remains accurately decodable. After this initial stage, all modules, including the U-3DGS components and auxiliary encoders, are jointly fine-tuned to enhance task-specific performance while preserving reconstruction fidelity.
Setup. Following the evaluation protocol established by SceneGraphLoc [13], we sample 123 distinct scenes from 30 rooms within the 3RScan test split. For each query image, the objective is to identify the correct scene from a candidate pool of 10 scenes, which includes the ground-truth scene. This experimental setup results in a total of 30,462 query evaluations.
Evaluation. At test time, each posed RGB image is encoded into a set of patch-level embeddings. For every candidate scene in the pool, these image patch embeddings are compared against all available object embeddings from that scene using cosine similarity. The final prediction for the scene is determined via a robust voting mechanism that aggregates all patch-object similarity scores. We report Recall@K for $K\in\{1,3,5\}$, which measures the frequency with which the correct scene appears among the top-K predicted scenes. This metric directly reflects the model’s capability to localize images effectively using the learned, object-centric multimodal representation.
Results. We compare our results with SceneGraphLoc [13] and the recent CrossOver method [22]. Our approach demonstrates competitive localization accuracy while crucially maintaining compatibility with 3D reconstruction and other downstream applications, a benefit stemming from our jointly trained, modular representation. Detailed results are presented in Table 4. In this table, we indicate whether a given method utilizes point clouds ($\mathcal{P}$), images ($\mathcal{I}$), other modalities such as object attributes and relationships ($\mathcal{O}$), or the proposed U-3DGS embedding. Our method, leveraging U-3DGS embeddings in conjunction with other modalities, achieves the highest Recall@1 and Recall@3 scores. This outcome suggests that the proposed U-3DGS embeddings furnish information comparable to, or even richer than, that provided by raw point clouds or images for the task of visual localization.
Furthermore, we present an ablation study for our method (indicated with an asterisk * in Table 4) using only the U-3DGS embeddings, without any specific fine-tuning of our main encoder for this localization task. As anticipated, the auxiliary modalities (attributes, relationships, etc.) offer valuable complementary information, contributing significantly to the superior performance of the full model. Interestingly, even in this constrained setting (U-3DGS embeddings alone, without targeted training), our method performs comparably to the recent CrossOver approach [22]. This highlights the inherent richness and suitability of our learned U-3DGS embeddings for visual localization tasks, even without explicit optimization for this specific application.
Method Modalities 10 scenes
[…]
SGLoc [13] 53.6 81.9 92.8
CrossOver [22] 46.0 77.9 90.5
Object-X 56.6 82.2 91.8
[…]
Object-X* 44.8 72.6 85.7
Table 4: Coarse visual localization on the 3RScan dataset [26] using the proposed U-3DGS embedding, compared to SceneGraphLoc [13] and CrossOver [22]. We report retrieval recall at 1, 3, and 5 when selecting the correct scene from 10 candidates. Evaluations are conducted using different map modalities: point cloud ($\mathcal{P}$), image ($\mathcal{I}$), other modalities ($\mathcal{O}$), and 3DGS. In the lower section (*), we also present results where the U-3DGS embedding is used without task-specific training.
Appendix D Supplementary: Scene Alignment
We provide additional details on the setup and evaluation procedure for the 3D Scene Alignment task on the 3RScan dataset [26], following the protocol established by SGAligner [21].
Setup. To construct the evaluation data, we generate sub-scenes by selecting fixed-length sequences of consecutive RGB-D frames from the 3RScan validation set. Each such sequence is then fused into a partial 3D reconstruction using volumetric integration. This process results in a total of 848 sub-scenes, each representing a distinct viewpoint or region within an original, larger scene. From these sub-scenes, we create 1,906 pairs by selecting pairs that originate from the same ground-truth scene. These pairs are deliberately constructed to span a wide range of spatial overlap percentages (from 10% to 90%) thereby ensuring coverage of both straightforward and challenging alignment scenarios.
Evaluation. We extract object embeddings independently from each sub-scene. These embeddings are produced by the same network architecture and weights trained for the scene localization task, as detailed in Section C. During the evaluation phase, we compute the cosine similarity between every object embedding in one sub-scene and all object embeddings in its paired sub-scene. For each object in the first sub-scene, candidate objects from the second sub-scene are ranked based on this similarity score. We evaluate the quality of these rankings using standard retrieval metrics: Mean Reciprocal Rank (MRR) and Hits@K, where $K\in\{1,2,\dots,5\}$. The Hits@K metric measures the proportion of queries for which a correct match appears within the top-K ranked results, while MRR quantifies the average inverse rank of the first correct match.
Appendix E Ablation Studies
This section analyzes the impact of key components and design choices in our proposed method. Table 5 presents results as a function of the compression rate, which is defined by the resolution of the underlying voxel grid, where each voxel stores eight parameters. As a reference, we also report results for standard 3D Gaussian Splatting (3DGS). In addition to our proposed 3D U-Net architecture for compression, we evaluate a naive downsampling approach that applies max pooling followed by interpolation.
[…]
Table 6 evaluates the robustness of our method to varying degrees of occlusion by systematically removing parts of an object before it is encoded. Occlusion is simulated by selecting a random point on the object’s surface and removing all geometry within a sphere of diameter $d$. The diameter $d$ is defined as a fraction of the object’s characteristic size; for example, $d=0.4$ corresponds to approximately 40% of the object’s volume being removed. The results indicate that even under severe occlusion, our proposed method maintains high reconstruction accuracy, thereby demonstrating its resilience to incomplete or missing input data.
Resolution Method LPIPS (Mean ± $\sigma$) $\downarrow$ Median $\downarrow$ PSNR (Mean ± $\sigma$) $\uparrow$ Median $\uparrow$
3DGS - $0.086\pm 0.082$ $0.060$ $30.15\pm 5.06$ $30.14$
[…]
$32^{3}$ Naive $0.124\pm 0.126$ $0.076$ $25.28\pm 5.51$ $25.80$
3D Unet $\mathbf{0.099\pm 0.108}$ $\mathbf{0.060}$ $\mathbf{27.06\pm 6.18}$ $\mathbf{26.84}$
$16^{3}$ Naive $0.189\pm 0.137$ $0.137$ $21.32\pm 5.53$ $21.41$
3D Unet $\mathbf{0.103\pm 0.113}$ $\mathbf{0.062}$ $\mathbf{27.01\pm 6.29}$ $\mathbf{26.71}$
$8^{3}$ Naive $0.257\pm 0.187$ $0.211$ $17.46\pm 5.26$ $16.86$
3D Unet $\mathbf{0.110\pm 0.119}$ $\mathbf{0.065}$ $\mathbf{26.74\pm 6.35}$ $\mathbf{26.50}$
Table 5: Ablation study on latent dimensions. Mean and median LPIPS and PSNR on a subset of scans from the test set. We compare the standard 3DGS (as a reference), the SLat embedding without dimensionality reduction, and U-3DGS with compressed representations at $32^{3}$, $16^{3}$, and $8^{3}$. Also, we evaluate naive downscaling approaches using max pooling and interpolation alongside the proposed 3D U-Net. The $16^{3}$ resolution is selected for all other experiments as it significantly reduces storage while maintaining near-optimal reconstruction accuracy.
$d$ LPIPS (Mean ± $\sigma$) $\downarrow$ Median $\downarrow$ PSNR (Mean ± $\sigma$) $\uparrow$ Median $\uparrow$
0.0 $0.104\pm 0.114$ $0.062$ $26.85\pm 6.25$ $26.66$
0.1 $0.104\pm 0.113$ $0.063$ $26.96\pm 6.32$ $26.69$
0.2 $0.106\pm 0.114$ $0.064$ $26.70\pm 6.44$ $26.50$
0.4 $0.113\pm 0.119$ $0.068$ $26.07\pm 6.70$ $25.93$
Table 6: Ablation study on occlusion. Before encoding an object, we randomly select a point on its surface and remove all parts within a spherical region of diameter $d$. For example, $d=0.4$ corresponds to a removal region spanning 40% of the object’s size. We report LPIPS and PSNR scores for different values of $d$ to assess the impact of occlusion on reconstruction quality.

### v2

@@
 The learned embedding enables 3D Gaussian Splatting-based object reconstruction, while also supporting a range of downstream tasks, including scene alignment, single-image 3D object reconstruction, and localization.
-Evaluations on two challenging real-world datasets demonstrate that Object-X produces high-fidelity novel-view synthesis comparable to standard 3D Gaussian Splatting, while significantly improving geometric accuracy.
+Evaluations on two challenging real-world datasets demonstrate that Object-X achieves high-fidelity novel-view synthesis comparable to standard 3D Gaussian Splatting, while significantly improving geometric accuracy.
 Moreover, Object-X achieves competitive performance with specialized methods in scene alignment and localization.
@@
 The embeddings learned from multi-modal data (e.g., mesh, images, text descriptions) enable fast 3D Gaussian Splat reconstruction via a specifically trained decoder, and other downstream tasks operating directly in the latent space, such as localization and scene alignment.
-Object-X allows for representing the scene as a set of object descriptors without having to store storage-heavy representations like point clouds and image databases, while providing similiar functionalities.
+Object-X allows for representing the scene as a set of object descriptors without having to store storage-heavy representations like point clouds and image databases, while providing similar functionalities.
 1 Introduction
@@
 Other modalities will be discussed when training also for downstream applications in the next section.
+We denote the resulting latent object embedding as $\mathbf{w}$ which is the central representation used throughout the paper.
 Voxelization and Feature Extraction.
@@
 Concatenating over all active voxels forms a sparse, voxel-aligned feature set as $f\;=\;\{(f_{i},p_{i})\}_{i=1}^{L},L\ll N^{3}$.
-Structured Latent Representation (SLat). We convert $f$ into a structured latent representation $z=\{(z_{i},p_{i})\}_{i=1}^{L}$ via a 3D encoder $E$ as follows:
+Structured Latent Representation (SLat). We convert $f$ into a structured latent representation $z=\{(z_{i},p_{i})\}_{i=1}^{L}$ via a 3D encoder $\mathcal{E}$ as follows:
 $z=\{(z_{i},p_{i})\}_{i=1}^{L}\;=\;\mathcal{E}(f),\;z_{i}\in\mathbb{R}^{C},\;p_{i}\in\{0,\dots,N-1\}^{3}.$
@@
 3.2 Compression to a Dense Embedding
-Although $z$ is sparse relative to raw 3D data, it can still be large for high-resolution grids when we aim at capturing fine details.
+Although $z$ is sparse relative to raw 3D data, it can still be large at high resolutions.
+Capturing fine details requires dense grids, which remain memory intensive.
 Moreover, downstream tasks favor a compact embedding for each object, rather than a collection of per-voxel vectors.
@@
 We then encourage $\hat{z}_{i}\approx z_{i}$ under a masked mean-square-error (MSE) or L1 loss that focuses on occupied voxels as $\hat{z}\;=\;f_{\text{decomp}}(\mathbf{w})$, and the objective includes
-$\mathcal{L}_{\text{compress}}\;=\;\frac{1}{L}\sum_{i=1}^{L}\Bigl{[}M_{i}\,\lVert\hat{z}_{i}-z_{i}\rVert^{2}\;+\;\tfrac{1-M_{i}}{w}\lVert\hat{z}_{i}-z_{i}\rVert^{2}\Bigr{]},$
+$\mathcal{L}_{\text{compress}}\;=\;\frac{1}{N^{3}}\sum_{i=1}^{N^{3}}\Bigl[M_{i}\,\lVert\hat{z}_{i}-z_{i}\rVert^{2}\;+\;\tfrac{1-M_{i}}{w}\lVert\hat{z}_{i}-z_{i}\rVert^{2}\Bigr],$
 where binary mask $M_{i}$ indicates if voxel $p_{i}$ is occupied, and $w$ is a down-weighting factor for non-occupied regions.
@@
 We next train a decoder $D_{\text{GS}}$ that takes in the reconstructed structured latents $\hat{z}$ (obtained from $\mathbf{w}$) and outputs a set of 3D Gaussian Splat parameters as follows:
-$\Theta\;=\;\mathcal{D}_{\text{GS}}(\hat{z}),\quad\Theta\;=\;\Bigl{\{}\!(x_{i},s_{i},q_{i},\alpha_{i},c_{i})\!\Bigr{\}}_{i=1}^{N}.$ (1)
+$\Theta\;=\;\mathcal{D}_{\text{GS}}(\hat{z}),\quad\Theta\;=\;\Bigl\{\!(x_{i},s_{i},q_{i},\alpha_{i},c_{i})\!\Bigr\}_{i=1}^{M}.$ (1)
 Each Gaussian is specified by position $x_{i}$, scale $s_{i}$, rotation $q_{i}$, opacity $\alpha_{i}$, and color $c_{i}$.
@@
 The optimization jointly minimizes the reconstruction loss, ensuring accurate recovery of the SLat representation from the U-3DGS embedding, and photometric loss, learning the mapping from SLat features into 3DGS that can be used for NVS and surface reconstruction.
-Figure 3: The proposed Object-X learns per-object embeddings that are beneficial for a number of downstream tasks, besides object-wise 3DGS reconstruction, such as cross-modal visual localization [13] (via image-to-object matching), 3D scene aligment [21] (via object-to-object matching), and full-scene reconstruction by integrating per-object Gaussians primitives.
+Figure 3: The proposed Object-X learns per-object embeddings that are beneficial for a number of downstream tasks, besides object-wise 3DGS reconstruction, such as cross-modal visual localization [14] (via image-to-object matching), 3D scene alignment [22] (via object-to-object matching), and full-scene reconstruction by integrating per-object Gaussians primitives.
 3.4 Learning Auxiliary Tasks
@@
 The newly introduced encoders for the auxiliary modalities are trained.
-In this stage, the learning is guided by a combined loss:
+In this stage, learning is guided only by the task-specific loss $\mathcal{L}_{\text{task}}(\mathbf{w}_{\text{concat}})$, where $\mathbf{w}_{\text{concat}}$ is the full concatenated embedding (U-3DGS + auxiliary features).
+For example, $\mathcal{L}_{\text{task}}$ may be a contrastive loss for localization.
+This ensures that the new auxiliary features are learned in a way that remains compatible with, and does not corrupt, the reconstructive capabilities of the core U-3DGS representation.
+Joint Fine-tuning: After the auxiliary encoders have been trained, all network components are unfrozen. The entire ensemble is then fine-tuned end-to-end using a combined objective:
 $\mathcal{L}_{\text{aux}}=\mathcal{L}_{\text{task}}(\mathbf{w}_{\text{concat}})+\lambda_{\text{recon}}\mathcal{L}_{\text{recon}}(\mathbf{w}_{\text{U-3DGS}}),$ (2)
-where $\mathbf{w}_{\text{concat}}$ is the full concatenated embedding (U-3DGS + auxiliary features).
-$\mathcal{L}_{\text{task}}$ is the task-specific loss (e.g., a contrastive loss for localization).
-$\mathcal{L}_{\text{recon}}$ is the reconstruction loss also used for pre-training the U-3DGS embeddings.
-Crucially, $\mathcal{L}_{\text{recon}}$ is applied using the U-3DGS decoder, which operates only on the corresponding U-3DGS portion ($\mathbf{w}_{\text{U-3DGS}}$) of $\mathbf{w}_{\text{concat}}$.
-This ensures that the new auxiliary features are learned in a way that remains compatible with, and does not corrupt, the reconstructive capabilities of the core U-3DGS representation.
+where $\mathcal{L}_{\text{recon}}$ is the reconstruction loss also used for pre-training the U-3DGS embeddings, applied using the U-3DGS decoder on the corresponding $\mathbf{w}_{\text{U-3DGS}}$.
 $\lambda_{\text{recon}}$ balances these two objectives.
-Joint Fine-tuning: After the auxiliary encoders have been trained, all network components are unfrozen.
-This includes the U-3DGS encoder and decoder, as well as all auxiliary modality encoders.
-The entire ensemble is then fine-tuned end-to-end using the same combined loss $\mathcal{L}_{\text{aux}}$ (Eq 2).
+This stage allows for mutual adaptation of all parts of the representation, further optimizing for both the specific auxiliary task and the foundational 3D reconstruction quality.
 Through this process, Object-X learns to effectively fuse intrinsic object properties (geometry, appearance via U-3DGS) with extrinsic or contextual information (text, relationships via auxiliary encoders).
 We will demonstrate this approach by training our model for visual localization [14], and subsequently evaluate its zero-shot or fine-tuned performance on related tasks like 3D scene alignment, showcasing the versatility and robustness of the learned augmented embeddings.
+4 Experiments
+Next, we will provide experiments on various tasks benefiting from Object-X. Ablation studies, more visuals, and detailed descriptions of baselines are provided in the supplementary material.
 Figure 4: Object reconstructions.
@@
 For each method, we present a rendered image from the reconstructed 3D Gaussians and the corresponding mesh.
-4 Experiments
-Next, we will provide experiments on various tasks benefiting from Object-X. Ablation studies, more visuals, and detailed descriptions of baselines are provided in the supplementary material.
 Mesh extraction.
@@
 Since ScanNet does not provide scene graph annotations, we apply SceneGraphFusion [30] on RGB-D sequences to generate 3D instance segmentations and object relationships (used for auxiliary tasks).
+This allows us to assess robustness to errors in 3D instance segmentation in the practical setting.
 Compared to 3RScan, ScanNet captures RGB-D sequences at a higher frame rate with minimal motion between consecutive frames.
@@
 3RScan 3DGS [28] 0.956 34.009 0.051 33.75 73.38 41.41 58.461 32.74
-[…]
+3DGS (12V) [28] 0.944 31.613 0.072 33.72 77.81 44.35 58.461 16.71
 2DGS (12V) [7] 0.932 29.613 0.093 26.55 67.69 35.67 84.280 16.71
@@
 ScanNet 3DGS [28] 0.975 38.138 0.032 33.81 79.13 43.00 129.02 270.97
-[…]
+3DGS (4V) [28] 0.949 30.754 0.059 36.95 82.37 48.18 129.02 155.26
 2DGS (4V) [7] 0.945 29.604 0.073 26.96 74.85 37.10 169.65 155.26
@@
 However, we also evaluate on ScanNet, where neither DepthSplat nor our method has been trained.
-For the baselines, we select $k$ frames that maximize viewpoint diversity.
-To achieve this, we apply $k$-means clustering to the positions of the cameras observing the object.
-In 3RScan, we use the maximum number of frames supported by DepthSplat (i.e., 12), while in ScanNet, where fewer non-overlapping frames are available, we limit the selection to four frames per object.
+For the baselines, we select $k$ frames that maximize viewpoint diversity by applying $k$-means clustering to the positions of the cameras observing the object.
+In 3RScan, we use the maximum number of frames supported by DepthSplat (i.e., 12), while in ScanNet, where fewer frames are available, we limit the selection to four views per object.
+To ensure fairness, 3DGS/2DGS are also evaluated under this sparse-view setting ($\leq 12$ views), matching the input regime of DepthSplat and reflecting the natural sparsity of 3RScan/ScanNet.
+DepthSplat itself is evaluated in its intended setup on unmasked scene-level images, with object masks applied only after reconstruction, ensuring a consistent and fair comparison.
 We present examples in Fig.
@@
 We omit geometric results for DepthSplat [32], which failed to produce reasonable geometry.
+We attribute the higher geometric accuracy to the voxel-grounded latent representation, which spatially constrains and aligns the decoded Gaussians to surfaces, leading to improved consistency compared to unstructured Gaussian sets.
 Our runtime is three orders of magnitude faster than methods relying on optimization.
@@
 Our geometric accuracy significantly outperforms all baselines.
-Given the higher resolution images in this dataset compared to 3RScan, the optimization-based approaches are quite slow, even when using only 4 views.
+Because ScanNet provides higher-resolution images than 3RScan, optimization-based approaches are considerably slower, even when using only 4 views.
 We are four orders of magnitude faster than 3DGS (4V) and 2DGS (4V) requiring $3$ ms to reconstruct an object on average.
@@
 Scene Reconstruction.
-Although the we do not explicitly perform scene reconstruction, we evaluate scene composition by integrating object-level reconstructions.
+Although we do not explicitly perform scene reconstruction, we evaluate scene composition by integrating object-level reconstructions.
 Object-X composes the full scene by decoding embeddings for each object and rendering their splats jointly.
@@
 Runtime remains comparable to the 12-view baselines, and significantly faster than full-scene 3DGS optimization.
+While we note that full-scene reconstruction is not the primary focus of this work, we include these experiments to demonstrate the viability of composing object-level embeddings into larger-scale reconstructions, where Object-X provides competitive quality, the highest geometric accuracy, and serves as a fast initializer for subsequent refinement.
 3.
@@
 We note that while there is still room for improvement in this task, achieving improved results demonstrate the versatility of the learned object embeddings.
+A promising direction, for example, is the principled integration of monocular 3D priors [33, 36, 29] could further enhance geometric fidelity without sacrificing efficiency.
+Additionally, we compare Object-X with the recent MIDI-3D [8], a diffusion-based generative model for single-image 3D reconstruction.
+Unlike Object-X, which encodes metrically grounded geometry from observed data, MIDI is a purely generative approach that learns to synthesize plausible 3D shapes directly from RGB inputs.
+The results are shown in Table 2.
+While MIDI achieves impressive visual quality on in-distribution images similar to its training data, its performance drops significantly on out-of-distribution scenes, such as those in 3RScan, where object shapes, scales, and textures differ from common internet imagery.
+In these cases, MIDI often produces geometrically inconsistent outputs that remain visually plausible but lack realism.
+In contrast, Object-X focuses on reconstruction rather than generation, leveraging voxel-grounded latent representations to maintain geometric consistency even under large appearance or domain shifts.
+As a result, it generalizes better to real-world scenes without requiring distribution-specific fine-tuning, producing reconstructions that are both structurally coherent and metrically faithful.
+Method Accuracy@0.05 $\uparrow$ Completion@0.05 $\uparrow$ F1@0.05 $\uparrow$
+MIDI [8] 29.522 44.342 34.516
+Object-X (RGB) 43.480 57.214 48.397
+Object-X (RGB-D) 65.780 61.759 63.223
+Table 2: Comparison with MIDI [8] on a subset of 3RScan. Metrics are computed at a 5 cm threshold. Methods marked with * use automatic alignment without manual registration.
 Table 3: Full-scene composition on 3RScan [27].
@@
 3RScan 3DGS [28] 0.855 24.404 0.417 16.48 31.99 21.40 260
-[…]
+3DGS (12V) [28] 0.767 18.806 0.517 16.81 32.63 21.88 150
 2DGS (12V) [7] 0.752 18.900 0.523 18.09 30.92 22.61 250
@@
 Method Modalities 10 scenes
-[…]
+$\mathcal{P}$ $\mathcal{I}$ O 3DGS R@1 R@3 R@5
 SGLoc [14] 53.6 81.9 92.8
@@
 In constrast to the baselines, Object-X is used without training on this task.
-Method Modalities Mean RR $\uparrow$ Hits@1 $\uparrow$ Hits@2 $\uparrow$ Hits@3 $\uparrow$ Hits@4 $\uparrow$ Hits@5 $\uparrow$
-$\mathcal{P}$ $\mathcal{O}$ 3DSG
+Method Modalities Mean $R$ $\uparrow$ Hits@1 $\uparrow$ Hits@2 $\uparrow$ Hits@3 $\uparrow$ Hits@4 $\uparrow$ Hits@5 $\uparrow$
+$\mathcal{P}$ $\mathcal{O}$ 3DGS
 EVA [12] 0.867 0.790 0.884 0.938 0.963 0.977
@@
 Our approach effectively bridges the gap between abstract learned object representations and detailed explicit 3D models, offering a scalable and practical solution for advanced 3D understanding.
-The code will be made public.
 Limitations.
@@
 The high degree of compression can lead to a loss of the finest details in some reconstructions, particularly for large or complex objects.
-Furthermore, while promising in zero-shot scenarios for tasks like single-image object reconstruction, performance may not yet consistently match optimized task-specific methods.
+Furthermore, while promising in zero-shot scenarios for tasks like single-image object reconstruction, performance does not yet consistently match that of optimized task-specific methods.
+Acknowledgements. This work was supported by an ETH Zurich Career Seed Award.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The abstract and introduction summarize the paper’s contributions.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9.
+Justification: Limitations are clearly discussed in the Conclusion section, highlighting challenges such as loss of fine detail in compressed representations and performance gaps in zero-shot tasks.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: The paper does not include theoretical results.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: Justification: The paper describes implementation details, model architecture, training setup, datasets used, and evaluation metrics in detail (see Section 4 and Implementation Details), additional implementation details will be included in the supplementary material.
+Furthermore, the code will be made public.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [No]
+  24. Justification: The code will be made public, but it is not yet available at submission time.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29.
+Justification: Detailed experimental settings, including datasets (3RScan, ScanNet), evaluation protocols, hardware used (A100 GPU), and hyperparameters, are provided in Section 4 and Implementation Details, additional implementation details will be included in the supplementary material.
+Furthermore code will be made public for further understanding.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [No]
+  34. Justification: The paper does not report error bars, confidence intervals, or significance tests. Results are shown as single point metrics (e.g., SSIM, PSNR, F1).
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: Section 4 provides hardware details (A100 GPU, 80GB RAM), per-object runtime, and comparisons to baseline methods in terms of compute efficiency.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: The paper does not violate NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [N/A]
+  49.
+Justification: The paper does not include a Broader Impact section because it focuses on research in 3D object representation.
+The work does not involve user-facing systems, sensitive data, or applications with immediate societal or ethical implications.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The model and data used do not pose high misuse risks and are standard in 3D vision research.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59.
+Justification: The paper uses publicly available datasets (3RScan, ScanNet) and pretrained models (e.g., DINOv2), all of which are properly cited.
+Licensing terms for these assets (e.g., academic non-commercial use for ScanNet) will be included in the supplementary materia
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64.
+Justification: The paper introduces new test splits and annotations for 3RScan and ScanNet, which are not included in the submission but will be publicly released with documentation upon publication.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
 Supplementary Material
@@
   5. 5.
-    Ablation results on compression, occlusion robustness, and architectural variants (Section E)
+    Single image reconstruction experiment (Section E)
+  6. 6.
+    Ablation results on compression, occlusion robustness, and architectural variants (Section F)
 Appendix A Baselines and Visualizations
@@
   2. 2.
-    3DGS ($k$V): Employs $k$ pre-selected views that observe the target object.
+    3DGS (kkV): Employs $k$ pre-selected views that observe the target object.
   3. 3.
-    2DGS ($k$V): Also uses $k$ pre-selected views observing the target object.
+    2DGS (kkV): Also uses $k$ pre-selected views observing the target object.
 To ensure a fair comparison by providing identical starting conditions for all methods, we initialize Gaussian splats directly from ground-truth object meshes.
@@
 Specifically, it undergoes an additional 4,000 iterations of 3DGS optimization.
-To ensure stability during this fine-tuning process, all learning rates are reduced by a factor of 10×10$\times$10 × compared to the standard 3DGS settings.
+To ensure stability during this fine-tuning process, all learning rates are reduced by a factor of $10\times$ compared to the standard 3DGS settings.
 Visualizations of scene-level reconstructions are presented in Figure 8.
@@
 We also test our method with only RGB input with depth predicted by Metric3D [6] to generate the initial point cloud.
-Similarly to the scene reconstruction case, we use Object-X to provide an initial reconstruction which we further refine by applying an additional 1,000 iterations of 3DGS optimization, using learning rates reduced by a factor of 10×10$\times$10 × (consistent with the scene-level refinement).
+Similarly to the scene reconstruction case, we use Object-X to provide an initial reconstruction which we further refine by applying an additional 1,000 iterations of 3DGS optimization, using learning rates reduced by a factor of $10\times$ (consistent with the scene-level refinement).
 Visual results for this setting are presented in Figure 7.
@@
 Method Modalities 10 scenes
-[…]
+$\mathcal{P}$ $\mathcal{I}$ O 3DGS Recall@1 Recall@3 Recall@5
 SGLoc [14] 53.6 81.9 92.8
@@
 Object-X 56.6 82.2 91.8
-[…]
+Object-X* 28.7 58.5 76.5
 Object-X* 44.8 72.6 85.7
@@
 The Hits@K metric measures the proportion of queries for which a correct match appears within the top-K ranked results, while MRR quantifies the average inverse rank of the first correct match.
-Appendix E Ablation Studies
+Appendix E Supplementary: Single-Image Reconstruction
+We provide qualitative results for the comparison of MIDI and Object X. The scenes on which the experiment is tested on are made available in the code.
+Figure 9: Qualitative comparison between MIDI [8] and Object-X on 3RScan single-image inputs.
+MIDI often produces realistic but misaligned shapes, while Object-X yields accurate and structurally coherent reconstructions that preserve scale and geometry.
+Appendix F Ablation Studies
 This section analyzes the impact of key components and design choices in our proposed method.
@@
 In addition to our proposed 3D U-Net architecture for compression, we evaluate a naive downsampling approach that applies max pooling followed by interpolation.
-[…]
+The results corresponding to a $64^{3}$ voxel grid resolution effectively represent our Structured Latent (SLat) representation without any subsequent compression, as this directly matches the original voxel resolution described in the main paper.
+The ablation results demonstrate that employing naive downsampling leads to a significant degradation in accuracy as the resolution decreases.
+In contrast, our proposed 3D U-Net maintains high fidelity with only a marginal loss in accuracy, while substantially reducing the number of parameters required per object from $64^{3}\times 8=2\,097\,152$ to a mere $8^{3}\times 8=4\,096$.
+Based on this analysis, we adopt a resolution of $16^{3}$ for the compressed representation in all our main experiments.
 Table 7 evaluates the robustness of our method to varying degrees of occlusion by systematically removing parts of an object before it is encoded.
@@
 Resolution Method LPIPS (Mean ± $\sigma$) $\downarrow$ Median $\downarrow$ PSNR (Mean ± $\sigma$) $\uparrow$ Median $\uparrow$
-3DGS - $0.086\pm 0.082$ $0.060$ $30.15\pm 5.06$ $30.14$
-[…]
-$32^{3}$ Naive $0.124\pm 0.126$ $0.076$ $25.28\pm 5.51$ $25.80$
+3DGS - $086\pm 0.082 0.0$600.060 $30.15\pm 5.06$ $30.14$
+$64^{3}$ (SLat) - $094\pm 0.101 0.0$590.059 $27.30\pm 6.13$ $27.06$
+$32^{3}$ Naive $124\pm 0.126 0.0$760.076 $25.28\pm 5.51$ $25.80$
 3D Unet $\mathbf{0.099\pm 0.108}$ $\mathbf{0.060}$ $\mathbf{27.06\pm 6.18}$ $\mathbf{26.84}$
@@
 $d$ LPIPS (Mean ± $\sigma$) $\downarrow$ Median $\downarrow$ PSNR (Mean ± $\sigma$) $\uparrow$ Median $\uparrow$
-0.0 $0.104\pm 0.114$ $0.062$ $26.85\pm 6.25$ $26.66$
-0.1 $0.104\pm 0.113$ $0.063$ $26.96\pm 6.32$ $26.69$
-0.2 $0.106\pm 0.114$ $0.064$ $26.70\pm 6.44$ $26.50$
-0.4 $0.113\pm 0.119$ $0.068$ $26.07\pm 6.70$ $25.93$
+0.0 $104\pm 0.114 0.0$620.062 $26.85\pm 6.25$ $26.66$
+0.1 $104\pm 0.113 0.0$630.063 $\pm 6.32 26.6926.69$
+0.2 $106\pm 0.114 0.0$640.064 $26.70\pm 6.44$ $26.50$
+0.4 $113\pm 0.119 0.0$680.068 $26.07\pm 6.70$ $25.93$
 Table 7: Ablation study on occlusion.